Write an expression for the Planck's law of radiation?
1
Expert's answer
2021-09-03T15:31:47-0400

Planck's law of radiation is given as follows:


"B(\nu, T) = \frac{ 2 h \nu^3}{c^2} \frac{1}{e^\frac{h\nu}{k_\mathrm B T} - 1}"

where "B(\nu, T)" is the spectral radiance of a body for frequency "\nu"  at absolute temperature "T", "k_B" is the Boltzmann constant, "h" is the Planck constant, and "c" is the speed of light in the medium.